A football coach starts at the 50
-yard line (A) and walks to the 10-yard line (B). He turns around and walks to the 40-yard line (C). He turns around again and walks back to the 0-yard line (D). He turns around once more and walks back to the 30-yard line (E), thus completing his 10 minutes of pacing. The coaches average velocity is.....
1
Expert's answer
2020-11-23T10:30:22-0500

Distance is:

"(50-10)+(40-10)+(40-0)+(30-0)=40+30+40+30=120 yards"

Velocity:

"V=\frac{S}{t}=\frac{120}{10}=12 yards/min"